OpenStudy (anonymous):

Find the slope of the line that passes through (3, -4) and (-1, 2). A. 2/3 B. 3/2 C. -2/3 D. -3/2

OpenStudy (anonymous):

-3/2, I think. (-)4-2 = -6 3-(-)1 = 4 -6/4 = -1.5 -3/2 = -1.5 So, the answer -3/2